OpenStudy (anonymous):

x intercept -2/5 and y intercept 5 find point slope and slope intercet form? i have y-5=25/2(x-0) for point slope and y=25/2x+5 for slope intercept form is this correct?

OpenStudy (whpalmer4):

Your known points are (-2/5, 0) and (0,5) correct?

OpenStudy (anonymous):

yes

OpenStudy (whpalmer4):

If so, your slope is (5-0)/(0-(-2/5)) = 25/2. point slope form would be either y-5 = (25/2)(x-0) or y-0 = (25/2)(x-(-2/5)) -> y = (25/2)(x+2/5) slope intercept form would be y = (25/2)x + 5 Answers look correct, good job!

OpenStudy (anonymous):

thank you